### v23

# Adrien CHAN-HON-TONG ONERA Université Paris Saclay, 91140, France

Abstract
Strict linear feasibility or linear separation is usually tackled using efficient approximation/stochastic algorithms (that may even run in sub-linear times in expectation). However, today state of the art for solving exactly/deterministically such instances is to cast them as a linear programming instances. Inversely, this paper introduces a self-concordant perceptron algorithm which tackles directly strict linear feasibility with interior point paradigm. This algorithm has the same worse times complexity than state of the art linear programming algorithms but it complexity can be characterized more precisely eventually proving that it binary complexity is low on a sub-family of linear feasibility.
1 Introduction
Linear separation is the problem of solving a system of linear inequality $Ax\geq\mathbf{1}$ where $A\in\mathbb{R}^{M\times N}$ is a given matrix and $x$ is a $N$ vector to find (assuming some exist). Trivially, this problem is equivalent to strict linear feasibility which aims to solve $Ax>\mathbf{0}$ as any solution of the strict linear feasibility can be up-scaled to become a solution of linear separation.
This problem can be found in many domains notably machine learning where it usually comes with additional regularization cost on $x$ leading to linear support vector machine (SVM) for binary classification: given $x_{1},...,x_{M}\in\mathbb{R}^{N}$ and $y_{1},...,y_{M}\in\{-1,1\}$, the hyperplane that maximizes the margin between positive and negative samples (assumed linearly separable) is given by
$\underset{w\in\mathbb{R}^{N},b\in\mathbb{R}}{\min}w^{T}w\ \ \ \mathrm{sb:}\ \forall m\in\{1,...,M\},\ y_{m}(w^{T}x_{m}+b)\geq 1$
Trivially, if one drops the $w^{T}w$ costs, the problem becomes to find $w,b$ such that $\forall m\in\{1,...,M\},\ y_{m}(w^{T}x_{m}+b)=y_{m}((w::b)^{T}(x_{m}::1))=(w::b)^{T}(y_{m}\times x_{m}::y_{m}))\geq 1$ which is a linear separation problem with $A$ the matrix $M\times(N+1)$ whose row $m$ is $y_{m}\times x_{m}::y_{m}$ where :::: stands for the concatenation.
This problem is usually approximated very efficiently in expectation. For example [13] proves that one can get a given approximation value with a number of iteration independent of $M$. Also, [7] offers a linear time approximation algorithm for linear SVM.
Now, for exact/deterministic solution, methods dedicated to linear feasibility like [5, 11] does not match the state of the art of linear programming. For example, [11] may require $O(NM^{2}\sqrt{M}\log(\frac{1}{\rho}))$ operations with $\rho$ being the maximal margin related to the instance i.e. $\rho=\underset{x\in\mathbb{R}^{N},x^{T}x\leq 1}{\max}\ \underset{m\in\{1,...,M\}}{\min}\ A_{m}x$ (in other words $\frac{1}{\rho}$ is the norm of the SVM related to $A$). Yet, if the matrix $A$ has total binary size $L$, this correspond to $O(NM^{2}\sqrt{M}L)$ in worse case, while state of the art of linear programming is currently $N^{\gamma}\sqrt{M}L$ operations (where $\gamma$ is the coefficient of matrix multiplication currently around 2.38) with [12, 10] and even $N^{\gamma-\frac{1}{2}}\sqrt{M}L$ with [14] a deterministic version of [4] which offers a data structure allowing to compress the time complexity of the sequence of matrix inversions (same number of iterations but with amortized lighter matrix inversions).
Thus, today, the best theoretical way to deal (exactly and deterministically) with a strict linear feasibility instance $Ax>\mathbf{0}$ is to cast it into a linear program like:
$\underset{x\in\mathbb{R}^{N},t\in\mathbb{R}}{\min}\ t\ \ \ \ \mathrm{sb:}\ Ax+t\mathbf{1}\geq\mathbf{0},\ \ -\mathbf{1}\leq x\leq\mathbf{1}$
This last problem is feasible with a trivial starting point $x=0,t=1$ and it is bounded with $t$ coarsely able to become $\rho$ (it would be exactly $\rho$ if the constraint was $x^{T}x\leq 1$ instead of $\mathbf{1}\leq x\leq\mathbf{1}$).
Then, this last problem can be solved classically by interior point method i.e. by considering
$G_{C}(x,t)=C\times t-\underset{m\in\{1,...,M\}}{\sum}\log(A_{m}x+t)$
[…]
Optimizing $x,t$ to reach $\underset{x,t}{\min}\ G_{C}(x,t)$ then increasing $C$ as when $C$ is large enough on can guarantee $t<0$.
Even more, if $L$ is the total binary size of $A$, one can prove that the calculus of this algorithm never require more than $O(L)$ bit for each variable, which is optimal in worse case.
2 Motivation
Despite binary property of state of the art linear programming algorithms are optimal in worse cases, no works have focused on characterizing the instances where binary size will be significantly lower than $L$. Indeed, even when convergence is very fast, binary size required at each step may be high as this algorithm (i.e. minimization of $G_{C}$) structurally increases $C$ precisely in order to allow the current point to go closer and closer to the boundary of the polyhedron $\{x,\ Ax+t\geq\mathbf{1},\ -\mathbf{1}\leq x\leq\mathbf{1}\}$. This makes it hard to bound precisely the binary dynamic of the variables through the resolution of the problem. In other words, the classical algorithms work by making the log converging toward a perfect barrier that would be 0 on strict positive values and infinity on 0 and negative ones, eventually leading some term to go very close to 0 even when encapsulated into a $\log$ and thus requiring careful rounding strategy.
Inversely, this paper offers an algorithm with same worse time complexity but whose binary times complexity can be precisely characterized: let $\mu=\underset{i,j}{\sum}|A_{i}A_{j}^{T}|$ and $\rho$ the margin (i.e. $\underset{x\in\mathbb{R}^{N},x^{T}x\leq 1}{\max}\ \underset{m\in\{1,...,M\}}{\min}\ A_{m}x$), then, the offered algorithm requires $O(\sqrt{M}\log(\frac{\sqrt{M}\mu}{\rho}))$ Newton steps with each variable having almost always a denominator bounded by $O(M)$ and a numerator bounded by $O(\log(\frac{\sqrt{M}\mu}{\rho}))$. In particular, for almost square ($N\approx M$) matrix densely filled with large number $O(\log(\mu))\approx\frac{L}{M^{2}}$ rather than $L$. If in addition the margin is high such that $\log(\frac{1}{\rho})\approx\frac{L}{M}$, then this algorithm is theoretically better by a factor $\frac{1}{M}$ for the binary size required during the algorithm than the conversion to a linear program where operations have not been proven possible with less than $L$ digits (as pointed table 1).
classical way offered algorithm
worse number of iteration $\sqrt{M}L$ $\sqrt{M}L$
precise number of iteration unclear $O(\sqrt{M}\log(\frac{\sqrt{M}\mu}{\rho}))$
worse number of bits/variable $L$ $L$
precise number of bits/variable unclear $O(\log(\frac{\sqrt{M}\mu}{\rho}))$
binary complexity on almost square unclear know to be
dense instance with high margin ($\leq O(M^{\gamma}\sqrt{M}L^{2})$) $O(M^{\gamma-2}\sqrt{M}L^{2})$
behavior of log influence decreases smoothly increase smoothly
known worse cases yes [1] not yet
Table 1: Complexity for solving a linear feasibility query $Ax>\mathbf{0}$ with $A\in\mathbb{R}^{M\times N}$ with total binary size $L$ by either solving the problem with state of the art linear program algorithms or with the self-concordant perceptron.
This algorithm has also two other interests:
  • The algorithm works by making the log terms more and more important rather then less and less important, this way it may not struggle on the same hard instances than classical path following algorithm which has currently be proven not strongly polynomial [1] on some hard instances. This may also be connected to the idea of converging to the center of the polyhedron rather than a vertex [6].
  • The proof is somehow simpler than the other of central-path based methods (no need to introduced neither the complementary condition nor the distance to the central path).
Let stress that the self-concordant perceptron described in this paper is a significantly improvement of the version from [3] where complexity was $O(ML)$ and not $O(\sqrt{M}L)$ thank to the adaptation of the path following process to this algorithm.
3 The self-concordant perceptron
3.1 Well-known theorems
This paper relies on the following well-known theorems.
3.1.1 Determinant bound:
If $A$ is a square matrix with total binary size $L$, then, $\log(Det(A))\leq L$. As a result, if $A$ is a matrix with total binary size $L$ (not necessarily square) and $\widetilde{A}$ is a square sub-matrix of $A$ with possibly $O(1)$ new rows and cols with $O(1)$ values, then, $\log(Det(\widetilde{A}))\leq L$. Combined with Cramer rules, it also implies that it $\chi$ verifies $\widetilde{A}x=\mathbf{1}$ (or any vector with $O(1)$ values), then, $\log(\chi^{T}\chi)\leq L$. This result is used to establish time complexity of linear programming algorithm since at least [8]. In particular, $\log\left(\underset{x,\ Ax\geq\mathbf{1}}{\min}x^{T}x\right)\leq O(L)$.
3.1.2 Self concordance theory:
If $\Psi(x)$ is a self-concordant function (mainly any sum of quadratic, linear, constant and $\log$ term), with a minimum $\Psi^{*}$, then, the Newton descent starting from $x_{start}$ allows to compute $x_{\epsilon}$ such as $\Psi(x_{\epsilon})-\Psi^{*}\leq\varepsilon$ in $\widetilde{O}(\Psi(x_{start})-\Psi^{*}+\log\log(\frac{1}{\varepsilon}))$ damped Newton steps. Each step is:
  • $\lambda_{\Psi}(x)\leftarrow\sqrt{(\nabla_{x}\Psi)^{T}(\nabla^{2}_{x}\Psi)^{-1}(\nabla_{x}\Psi)}$
  • $x\leftarrow x-\frac{1}{1+\lambda_{\Psi}(x)}(\nabla^{2}_{x}\Psi)^{-1}(\nabla_{x}\Psi)$
Importantly, $(x-x^{*})^{T}(\nabla^{2}_{x}\Psi)^{-1}(x-x^{*})\leq\frac{\varepsilon}{1-\varepsilon}$ when when $\Psi(x_{\epsilon})-\Psi^{*}\leq\varepsilon$ (for $\varepsilon$ below 1). The proof of this central theorem can be found in [9, 2].
3.2 The self concordant Perceptron
The self concordant Perceptron is described as a pseudo code in table 2.
def self-concordant-perceptron(A)(A) :
  1. 1.
    $\varepsilon\leftarrow\frac{1}{2^{O(L)}}$, $\delta\leftarrow 1$, $v\leftarrow\mathbf{1}$, $\mu\leftarrow\underset{i,j}{\sum}|A_{i}A^{T}_{j}|$
  2. 2.
    while $\neg(AA^{T}v>\mathbf{0})$
    1. (a)
      symbolically $F_{\delta}(u)=\delta\times\mathbf{1}^{T}u+\frac{u^{T}AA^{T}u}{\mu}-\underset{m\in\{1,...,M\}}{\sum}\log(u_{m})$
    2. (b)
      $v\leftarrow$ rounded Newton descent from $v$ on $F_{\delta}(u)$ with $\varepsilon$ precision
    3. (c)
      $\delta\leftarrow\delta\times(1-\frac{1}{\sqrt{M}})$
Table 2: Pseudo code of the self concordant perceptron.
More mathematically, given $A\in\mathbb{Z}^{M\times N}$ a feasibility instance with $\mu=\underset{i,j}{\sum}|A_{i}A_{j}^{T}|$, the algorithm is mainly a Newton descent on the function $\forall v,\ F_{\delta}(v)=\delta\times\mathbf{1}^{T}v+\frac{v^{T}AA^{T}v}{\mu}-\underset{m\in\{1,...,M\}}{\sum}\log(v_{m})$ to reach $\min_{v}F_{\delta}(v)$ then decaying $\delta$ to move into the central-path related to $F_{\delta}$.
To prove the claim presented in table 1, this paper will proceed as follow:
  • first, it will be proven that the central path converge toward a solution,
  • then, the complexity of the algorithm will be established from mathematical point of view (without rounding consideration),
  • finally, it will be proven that the rounding does not harm the process.
3.3 Convergence of the central path
3.3.1 Lemma on the lower bound of the quadratic term
Let $A\in\mathbb{Z}^{M\times N}$ with the assumption that there is a solution to $Ax\geq\mathbf{1}$ and let $x$ the SVM related to $A$ i.e. $x=\underset{\chi\in\mathbb{R}^{N}}{\arg\min}\ \chi^{T}\chi\ \ \ \ \mathrm{sb:}Ax\geq\mathbf{1}$.
[…]
proof
[…]
This inequality directly implies the second by injecting of $A^{T}v||_{2}^{2}\leq M\mu$.
3.3.2 Optimality conditions
$\forall\delta>0,v>\mathbf{0},v^{T}AA^{T}v\geq 0$ thus $F_{\delta}(v)=\delta\times\mathbf{1}^{T}v+\frac{v^{T}AA^{T}v}{\mu}-\underset{m\in\{1,...,M\}}{\sum}\log(v_{m})\geq\underset{m\in\{1,...,M\}}{\sum}\delta v_{m}-\log(v_{m})$. Thus, $F_{\delta}(v)$ is trivially lower bounded. As $F_{\delta}$ is convex, it has a minimum $w(\delta)=\underset{v\in\mathbb{R}^{M}}{\arg\min}\ F_{\delta}(v)$.
Now, from null gradient condition, it comes that
$\forall\delta>0,m\in\{1,...,M\},\delta+2\frac{A_{m}A^{T}w(\delta)}{\mu}-\frac{1}{w_{m}(\delta)}=0$
In particular, $\forall\delta>0,m\in\{1,...,M\},\delta\times w_{m}(\delta)+2w_{m}(\delta)\times\frac{A_{m}A^{T}w(\delta)}{\mu}=1$. By summing over $m$, $\forall\delta>0,\delta\times\mathbf{1}^{T}w(\delta)+2\frac{w(\delta)^{T}A_{m}A^{T}w(\delta)}{\mu}=M$ thus, $2w(\delta)^{T}AA^{T}w(\delta)<M\mu$ as $w>\mathbf{0}$ implies $\mathbf{1}^{T}w(\delta)>0$.
3.3.3 Solution
Combining lemma 3.3.1 and 3.3.2, it comes that $\forall\delta>0,m\in\{1,...,M\}$
$w_{m}(\delta)\leq\frac{\sqrt{M}\mu}{\rho},\mathrm{and},\ \delta+\frac{2A_{m}A^{T}w(\delta)}{\mu}-\frac{1}{w_{m}(\delta)}=0$
In particular, for $\forall\delta>0$
$\delta<\frac{\rho}{\sqrt{M}\mu}\Rightarrow\ \frac{2A_{m}A^{T}w(\delta)}{\mu}=\frac{1}{w_{m}(\delta)}-\delta>0$
$\forall 0<\delta<\frac{\rho}{\sqrt{M}\mu},\ AA^{T}w(\delta)>\mathbf{0}$
Importantly, the existence of $w(\delta)$ is trivial proven for any $\delta>0$, but the fact that $w$ converges toward a solution is only true when there exists a solution to the linear feasibility instance (i.e. $w(0)$ exists) as it requires the inequality 3.3.1: for large $\delta$ the linear term forbids $w$ to be too large but allows for not being a solution, while for small $\delta$, the quadratic term takes over the role of preventing $w$ to be large yet forcing it to converge toward a solution as $\delta$ is then negligible.
3.3.4 Number of decaying steps
As the algorithm stops at least when $\delta\leq\frac{\rho}{\sqrt{M}\mu}$, and that each step decay $\delta$ by a factor $(1-\frac{1}{\sqrt{M}})$, then, the number of steps verifies $(1-\frac{1}{\sqrt{M}})^{K}\leq\frac{\rho}{\sqrt{M}\mu}$ i.e. $k\leq\sqrt{M}\log(\frac{\sqrt{M}\mu}{\rho})$.
However, at this point step $c$ may require many Newton steps. But, this paper proves that in average (deterministically) it does not.
3.4 Complexity
3.4.1 Initialization
By definition, $F_{1}(v)=\mathbf{1}^{T}v+\frac{v^{T}AA^{T}v}{\mu}-\underset{m\in\{1,...,M\}}{\sum}\log(v_{m})\geq\underset{m\in\{1,...,M\}}{\sum}v_{m}-\log(v_{m})\geq M$ as the function $\psi(u)=u-\log(u)$ verifies $\psi^{\prime}(u)=1-\frac{1}{u}$ eventually $\phi(u)\geq\phi(1)=1$.
Independently, $F_{1}(\mathbf{1})\leq M+1$. Thus, it takes only $O(1+\log\log(\frac{1}{\varepsilon}))$ to get a $\varepsilon$-approximation of $w(1)$ from $\mathbf{1}$.
3.4.2 Centering cost
Let first remark that optimality condition 3.3.2 not only implies that $\frac{w(\delta)AA^{T}w(\delta)}{\mu}\leq M$ as $\delta\mathbf{1}^{T}w(\delta)>0$ but, also that $\delta\mathbf{1}^{T}w(\delta)<M$ as $w(\delta)AA^{T}w(\delta)>0$. Then
$F_{(1-\frac{1}{\sqrt{M}})\delta}(w(\delta))=F(w(\delta),\delta)+((1-\frac{1}{\sqrt{M}})\delta-\delta)\mathbf{1}^{T}w(\delta)$
$F(w(\delta),\delta)-\frac{\delta\mathbf{1}^{T}w(\delta)}{\sqrt{M}}$
But, $F_{\delta}(w((1-\frac{1}{\sqrt{M}})\delta)=F_{(1-\frac{1}{\sqrt{M}})\delta}(w((1-\frac{1}{\sqrt{M}})\delta)$
$(\delta-(1-\frac{1}{\sqrt{M}})\delta)\mathbf{1}^{T}w((1-\frac{1}{\sqrt{M}})\delta)$
$F_{(1-\frac{1}{\sqrt{M}})\delta}(w((1-\frac{1}{\sqrt{M}})\delta)$
$(\frac{1}{1-\frac{1}{\sqrt{M}}}-1)\times(1-\frac{1}{\sqrt{M}})\delta\mathbf{1}^{T}w((1-\frac{1}{\sqrt{M}})\delta)$
$F_{(1-\frac{1}{\sqrt{M}})\delta}(w((1-\frac{1}{\sqrt{M}})\delta)$
$(\frac{\sqrt{M}}{\sqrt{M}-1}-\frac{\sqrt{M}-1}{\sqrt{M}-1})\times(1-\frac{1}{\sqrt{M}})\delta\mathbf{1}^{T}w((1-\frac{1}{\sqrt{M}})\delta)$
$F_{(1-\frac{1}{\sqrt{M}})\delta}(w((1-\frac{1}{\sqrt{M}})\delta)$ $\frac{(1-\frac{1}{\sqrt{M}})\delta\mathbf{1}^{T}w((1-\frac{1}{\sqrt{M}})\delta)}{\sqrt{M}-1}$
By combining the two
$F_{(1-\frac{1}{\sqrt{M}})\delta}(w(\delta))=F_{(1-\frac{1}{\sqrt{M}})\delta}(w((1-\frac{1}{\sqrt{M}})\delta)$
$\frac{\delta\mathbf{1}^{T}w(\delta)}{\sqrt{M}}+\frac{(1-\frac{1}{\sqrt{M}})\delta\mathbf{1}^{T}w((1-\frac{1}{\sqrt{M}})\delta)}{\sqrt{M}-1}$
Thus, the cost of an individual step 2.b of algorithm table 1 is
$O\left(\frac{(1-\frac{1}{\sqrt{M}})\delta\mathbf{1}^{T}w((1-\frac{1}{\sqrt{M}})\delta)}{\sqrt{M}-1}-\frac{\delta\mathbf{1}^{T}w(\delta)}{\sqrt{M}}+\log\log(\frac{1}{\varepsilon})\right)$
Newton steps, and, this can be as large as $\sqrt{M}$ if $(1-\frac{1}{\sqrt{M}})\delta\mathbf{1}^{T}w((1-\frac{1}{\sqrt{M}})\delta)=O(M)$ and $\delta\mathbf{1}^{T}w(\delta)=O(1)$.
But, let first remark that if $(1-\frac{1}{\sqrt{M}})\delta\mathbf{1}^{T}w((1-\frac{1}{\sqrt{M}})\delta)\leq\delta\mathbf{1}^{T}w(\delta)$ (i.e. the quantity $\delta\mathbf{1}^{T}w(\delta)$ decreases after $\delta$ decay), then, such step only costs at most $O(1+\log\log(\frac{1}{\varepsilon}))$ as $M(\frac{1}{\sqrt{M}-1}-\frac{1}{\sqrt{M}})=\frac{M}{M-\sqrt{M}}\underset{M\rightarrow\infty}{\rightarrow}1$.
And, on the other hand, $F_{(1-\frac{1}{\sqrt{M}})\delta}(w(\delta))\leq F_{(1-\frac{1}{\sqrt{M}})\delta}(w((1-\frac{1}{\sqrt{M}})\delta)$. Thus,
$\frac{(1-\frac{1}{\sqrt{M}})\delta\mathbf{1}^{T}w((1-\frac{1}{\sqrt{M}})\delta)}{\sqrt{M}-1}-\frac{\delta\mathbf{1}^{T}w(\delta)}{\sqrt{M}}\geq 0$
$(1-\frac{1}{\sqrt{M}})\delta\mathbf{1}^{T}w((1-\frac{1}{\sqrt{M}})\delta)\geq(1-\frac{1}{\sqrt{M}})\times\delta\mathbf{1}^{T}w(\delta)\geq\delta\mathbf{1}^{T}w(\delta)-\sqrt{M}$
Thus, a $\sqrt{M}$-cost step (corresponding to the worse case $\delta\mathbf{1}^{T}w(\delta)=O(1)$ while $(1-\frac{1}{\sqrt{M}})\delta\mathbf{1}^{T}w((1-\frac{1}{\sqrt{M}})\delta)=O(M)$) can only happen after $\sqrt{M}$ decreasing steps from an hypothetical situation where $\delta\mathbf{1}^{T}w(\delta)=O(M)$.
So if one consider not only a single step but the set of all steps where large increase of $\delta\mathbf{1}^{T}w(\delta)$ between $\delta$ and $(1-\frac{1}{\sqrt{M}})\delta$ should necessarily be preceded by large number of steps with decreasing $\delta\mathbf{1}^{T}w(\delta)$, then the amortized cost of a step is $O(1+\log\log(\frac{1}{\varepsilon}))$.
In other words, for any given step where $\delta\mathbf{1}^{T}w(\delta)$ is decreasing, the algorithm should spend twice $O(1)$ Newton steps. One $O(1)$ Newton steps to go from $w(\delta)$ to $w((1-\frac{1}{\sqrt{M}})\delta)$, and, one $O(1)$ in the future to return from $(1-\frac{1}{\sqrt{M}})\delta\mathbf{1}^{T}w((1-\frac{1}{\sqrt{M}})\delta)$ level to $\delta\mathbf{1}^{T}w(\delta)$. Yet, it is still $O(1)$ when amortizing.
Importantly, it is not a $O(1)$ on expectation (which would be related to a randomized algorithm), it is $O(1)$ with one part that would be done after.
3.4.3 Mathematical complexity
As step 2.b costs $O(1+\log\log(\varepsilon))$ in average, then, the total complexity is $O\left(M^{\gamma}\sqrt{M}\log(\frac{\sqrt{M}\mu}{\rho})(1+\log\log(\frac{1}{\varepsilon}))\right)$ which is $O(M^{\gamma}\sqrt{M}L)$ in worse case, but which is better characterized.
In particular, on instance where $\log(\frac{\mu}{\rho})\ll L$, then, this algorithm is interesting.
Now, the classical way may also benefit from the fact that $\log(\frac{\mu}{\rho})\ll L$. Yet, this algorithm has an other interesting property: $v\leq\log(\frac{\mu}{\rho})$ and all log terms are large. Yet, this last property seems structurally hard for the classical way where log terms tend to 0.
Thus, complexity of this algorithm is equivalent to the one of the classical way, but with a better characterization which may however be shared with the classical way. But, the binary property of this algorithm may be unique in those last easy instances.
3.5 Effect of rounding
3.5.1 variable rounding
The interest of this algorithm is that ceiling $v$ during Newton step can be almost trivially. Indeed, $\underset{m}{\sum}\log(v_{m}+\zeta_{m})>\underset{m}{\sum}\log(v_{m})$ if $\zeta\geq\mathbf{0}$ Thus, to design a ceiling process, one only has to consider $\delta\mathbf{1}^{T}(v+\zeta)+\frac{(v+\zeta)^{T}AA^{T}(v+\zeta)}{\mu}$ with the additional knowledge that $\frac{v^{T}AA^{T}v}{\mu}\leq M$.
[…]
Thus, a simple periodic ceiling with denominator $O(M)$ is sufficient to maintain a $O(1)$ errors which can be compensated with $O(1)$ Newton steps.
3.5.2 weight rounding
In the same way $(\delta+\zeta)\mathbf{1}^{T}v$ can be harmless if $\zeta\leq\frac{\delta}{M}$ as in this case, $(\delta+\zeta)\mathbf{1}^{T}v-\delta\mathbf{1}^{T}v\leq\frac{\delta\mathbf{1}^{T}v}{M}\leq 1$.